OpenStudy (anonymous):

need to find the area of their front and side yards since they want to reseed the yard. Both side yards mearsure 3 meters by 10 meters while the front yard is a square with a side of 9 meters. They do not need to reseed the portion of the front yard covering 16 square meters where the flower bed is located. What is the area of the yard that the want to reseed?

OpenStudy (anonymous):

Side yard = 3*10 = 30m^2 Front yard = 9^2 =81m^2 Total area of yard is the two side yard areas plus the front yard area 2*30 + 81 = 141m^2 The area to be reseeded is simply 16m^2 less than the total yard area 141 - 16 = 125m^2